Solve. -3 (10x-3) + 10 (3x-3) = x-7

Answer by tutorcecilia(2152)

First, distribute the -3 and the 10:
-3 (10x-3) + 10 (3x-3) = x-7
-30x + 9 + 30x - 30 = x-7
Combine like-terms:
-30x + 30x + 9 - 30 = x - 7
Simplify:
-21 = x - 7
Solve for x
-21 + 7 = x - 7 + 7
-14 = x
.
Check by plugging x=-14 back into the original equation:
-3 (10(-14)-3) + 10 (3(-14)-3) = (-14)-7
Simplify and solve:
-3 (-140-3) + 10 (-42 - 3) = -21
-3 (-140-3) + 10 (-42 - 3) = -21
-3 (-143) + 10 (-45) = -21
-3 (-143) + 10 (-45) = -21
429 + (-450) = -21
-21 = -21 The answer checks out.
If you need additional one-on-one math help, feel free to e-mail your problems to tutorcecilia@yahoo.com